# Rank Two Sheaves With Low Discriminant on the Fano Threefold of Index 2 and Degree 5

### v1

@@ 4. Partial Results and a Conjecture
Let $E\in\operatorname{Coh}^{\beta}(X)$ be a tilt-semistable object with $\mathrm{ch}(E)=H+dH^{2}+eH^{3}$. Then $e\leq\frac{1}{24}+\frac{d^{2}}{2}.$
Proof.
@@
The proof of [17, Lemma 3.3] holds true for any smooth projective threefold over $\mathbb{C}$, for which the generalized Bogomolov-Gieseker inequality (Proposition 2.2) holds, in particular, for $X_{5}$.∎
Lemma 4.2.
@@
Let $E\in\operatorname{Coh}^{\beta}(X)$ be a tilt-semistable object with $\mathrm{ch}(E)=1-dH^{2}+eH^{3}$, then $e\leq\frac{d(d+1)}{2}$.
Proof.
@@
Note that $H^{2}\cdot\mathrm{ch}^{-1}_{1}(E)$ has the least positive value, so $Q_{0,-1}(E)\geq 0$. This, together with the fact that $e\in\frac{1}{10}\mathbb{Z}$, implies that for $d=0$ or $d=\frac{1}{5}$ we have $e\leq 0$; for $d=\frac{2}{5}$ we have $e\leq\frac{1}{5}$; for $d=\frac{3}{5}$ we have $e\leq\frac{2}{5}$ and for $d=\frac{4}{5}$ we have $e\leq\frac{3}{5}$. For $d\geq 1$ the statement follows from arguments in the proof of [17, Proposition 3.2].∎
@@
Denote by $S$ a smooth hyperplane section of $X\subset\mathbb{P}^{6}$. It is known that $S$ is a del Pezzo surface of degree 5, isomorphic to a blow up of $\mathbb{P}^{2}$ in 4 generic points. Denote by $p:S\to\mathbb{P}^{2}$ the projection of the blow up, denote by $H^{\prime}$ the class of $p^{*}(H)$ in $\mathrm{Cl}\ S$ and by $E_{1},E_{2},E_{3},E_{4}$ the exceptional divisors corresponding to the blown up points. We can compute that $c_{1}(T_{S})=-c_{1}(\omega_{S})=3H^{\prime}-\sum_{i=1}^{4}E_{i}$ and $c_{2}(T_{S})=c_{2}(T_{\mathbb{P}^{2}})+4=7$ (using [10, Example 15.4.3]), which implies that $\mathrm{td}(T_{S})=1+\frac{3H^{\prime}-\sum_{i=1}^{4}E_{i}}{2}+P$ (here $P$ is the class of a point). Denote by $D$ the divisor $aH^{\prime}+\sum_{i=1}^{4}b_{i}E_{i}$ and by $i:S\to X$ the canonical injection.
Lemma 4.3.
@@
Proof.
Note that the embedding $S\to\mathbb{P}^{5}$ is given by the anticanonical linear system $|3H^{\prime}-\sum_{i=1}^{4}E_{i}|$ (indeed, adjunction formula implies that $\omega_{S}^{\vee}\cong\mathcal{O}_{X}(1)|_{S}$). It follows that $i_{*}(H^{\prime})=(3H^{\prime}-\sum_{i=1}^{4}E_{i},H^{\prime})L=\frac{3H^{2}}{5}$ and $i_{*}(E_{j})=(3H^{\prime}-\sum_{i=1}^{4}E_{i},E_{j})L=\frac{H^{2}}{5}$ for all $j$. Now a straightforward calculation using the Grothendieck–Riemann–Roch Theorem gives the statement of the lemma.∎
Proposition 4.4.
@@
Proof.
Lemma 4.3 implies that for a fixed $C$ maximizing $\mathrm{ch}_{3}(i_{*}\mathcal{O}_{S}(D))$ is equivalent to maximizing $Q:=a^{2}-\sum_{i=1}^{4}b_{i}^{2}$. For a fixed $C=3a+\sum_{i=1}^{4}b_{i}$, a fixed $a\in\mathbb{R}$ and varying $(b_{i})\in\mathbb{R}^{4}$ the maximal value of $Q$ is attained when all $b_{i}$ are equal to $\frac{C-3a}{4}$. The values of $b_{i}$ in Proposition 4.4 are obtained by choosing a point $\textbf{b}=(b_{i})\in\mathbb{Z}^{4}$ with $\sum_{i=1}^{4}b_{i}=C-3a$ that is closest to the point with all $b_{i}$ equal to $\frac{C-3a}{4}$ in the usual Euclidean metric. By Pythagoras’ Theorem this b is indeed a point minimizing $\sum_{i=1}^{4}b_{i}^{2}$ and maximizing $Q$ for fixed $C$ and $a$. It remains to prove that maximum of $Q$ can be reached only for the values of $a$ given in the Proposition.
For fixed $C$ and $a$ we have $Q\leq f_{C,a}:=a^{2}-4\left(\frac{C-3a}{4}\right)^{2}=-\frac{5}{4}a^{2}+\frac{3}{2}Ca-\frac{C^{2}}{4}$. This is a parabola in $a$ with a maximum at $a_{0}=\frac{3C}{5}$. This proves the Proposition in the case $C=5k,k\in\mathbb{Z}$. In the case $C=5k+1,k\in\mathbb{Z}$ $a_{0}=\frac{3C}{5}$ lies in the interval $(3k,3k+1)$. For the values of $a,b_{i}$ from Proposition 4.4 (2) we have $Q=5k^{2}+2k-1$. A calculation gives $f_{5k+1,3k-1}=5k^{2}+2k-3$. Hence for all $a\leq 3k-1$ $Q\leq f_{5k+1,a}\leq f_{5k+1,3k-1}<5k^{2}+2k-1$. Also we find $f_{5k+1,3k+2}=5k^{2}+2k-\frac{9}{4}$, hence for all $a\geq 3k+2$ $Q\leq 5k^{2}+2k-\frac{9}{4}<5k^{2}+2k-1$. Hence $a\in\{3k,3k+1\}$. This proves the case (2). The cases (3-5) can be proven completely analogously. ∎
@@
    (19) $0\to\mathcal{O}_{X}(-1)^{\oplus 2}\to E\to\mathcal{O}_{S}(D)\to 0.$
  2. (2)
@@
Proof.
We can calculate that the radius $\rho_{Q}$ of the numerical wall $Q_{\alpha,\beta}(E)=0$ satisfies the inequality
@@
Bounding each term, we see that $\frac{\rho_{Q}}{\overline{\Delta}_{H}(E)}>\frac{1}{300}$, hence we can apply Proposition 2.4 (1) to get the statement of our lemma.∎
Lemma 4.7.
@@
Proof.
Follows along the same lines as the proof of [23, Lemma 3.9]. Note that in our case
@@
Since the step by which $\mathrm{ch}_{3}(E)$ can change for a fixed $\mathrm{ch}_{\leq 2}(E)$ is equal to $\frac{H^{3}}{10}$, the bound proven in [23, Lemma 3.9] suffices to prove our lemma.∎

### v2

@@ 4 Partial Results and a Conjecture
Let $E\in\mathrm{Coh}^{\beta}(X)$ be a tilt-semistable object with $\mathrm{ch}(E)=H+dH^{2}+eH^{3}$. Then $e\leq\frac{1}{24}+\frac{d^{2}}{2}.$
@@
Proof. The proof of [17, Lemma 3.3] holds true for any smooth projective threefold over $\mathbb{C}$, for which the generalized Bogomolov-Gieseker inequality (Proposition 2.2) holds, in particular, for $X_{5}$. $\square$
Lemma 4.2
@@
Let $E\in\mathrm{Coh}^{\beta}(X)$ be a tilt-semistable object with $\mathrm{ch}(E)=1-dH^{2}+eH^{3}$, then $e\leq\frac{d(d+1)}{2}$.
@@
Proof. Note that $H^{2}\cdot\mathrm{ch}^{-1}_{1}(E)$ has the least positive value, so $Q_{0,-1}(E)\geq 0$. This, together with the fact that $e\in\frac{1}{10}\mathbb{Z}$, implies that for $d=0$ or $d=\frac{1}{5}$ we have $e\leq 0$; for $d=\frac{2}{5}$ we have $e\leq\frac{1}{5}$; for $d=\frac{3}{5}$ we have $e\leq\frac{2}{5}$ and for $d=\frac{4}{5}$ we have $e\leq\frac{3}{5}$. For $d\geq 1$ the statement follows from arguments in the proof of [17, Proposition 3.2]. $\square$
@@
Denote by $S$ a smooth hyperplane section of $X\subset\mathbb{P}^{6}$. It is known that $S$ is a del Pezzo surface of degree 5, isomorphic to a blow up of $\mathbb{P}^{2}$ in 4 generic points. Denote by $p:S\to\mathbb{P}^{2}$ the projection of the blow up, denote by $H^{\prime}$ the class of $p^{*}(H)$ in $\mathrm{Cl}\ S$ and by $E_{1},E_{2},E_{3},E_{4}$ the exceptional divisors corresponding to the blown up points. We can compute that $c_{1}(T_{S})=-c_{1}(\omega_{S})=3H^{\prime}-\sum_{i=1}^{4}E_{i}$ and $c_{2}(T_{S})=c_{2}(T_{\mathbb{P}^{2}})+4=7$ (using [9, Example 15.4.3]), which implies that $\mathrm{td}(T_{S})=1+\frac{3H^{\prime}-\sum_{i=1}^{4}E_{i}}{2}+P$ (here $P$ is the class of a point). Denote by $D$ the divisor $aH^{\prime}+\sum_{i=1}^{4}b_{i}E_{i}$ and by $i:S\to X$ the canonical injection.
Lemma 4.3
@@
Proof. Note that the embedding $S\to\mathbb{P}^{5}$ is given by the anticanonical linear system $|3H^{\prime}-\sum_{i=1}^{4}E_{i}|$ (indeed, adjunction formula implies that $\omega_{S}^{\vee}\cong\mathcal{O}_{X}(1)|_{S}$). It follows that $i_{*}(H^{\prime})=(3H^{\prime}-\sum_{i=1}^{4}E_{i},H^{\prime})L=\frac{3H^{2}}{5}$ and $i_{*}(E_{j})=(3H^{\prime}-\sum_{i=1}^{4}E_{i},E_{j})L=\frac{H^{2}}{5}$ for all $j$. Now a straightforward calculation using the Grothendieck–Riemann–Roch Theorem gives the statement of the lemma. $\square$
Proposition 4.4
@@
Proof. Lemma 4.3 implies that for a fixed $C$ maximizing $\mathrm{ch}_{3}(i_{*}\mathcal{O}_{S}(D))$ is equivalent to maximizing $Q:=a^{2}-\sum_{i=1}^{4}b_{i}^{2}$. For a fixed $C=3a+\sum_{i=1}^{4}b_{i}$, a fixed $a\in\mathbb{R}$ and varying $(b_{i})\in\mathbb{R}^{4}$ the maximal value of $Q$ is attained when all $b_{i}$ are equal to $\frac{C-3a}{4}$. The values of $b_{i}$ in Proposition 4.4 are obtained by choosing a point $\textbf{b}=(b_{i})\in\mathbb{Z}^{4}$ with $\sum_{i=1}^{4}b_{i}=C-3a$ that is closest to the point with all $b_{i}$ equal to $\frac{C-3a}{4}$ in the usual Euclidean metric. By Pythagoras’ Theorem this b is indeed a point minimizing $\sum_{i=1}^{4}b_{i}^{2}$ and maximizing $Q$ for fixed $C$ and $a$. It remains to prove that maximum of $Q$ can be reached only for the values of $a$ given in the Proposition.
For fixed $C$ and $a$ we have $Q\leq f_{C,a}:=a^{2}-4\left(\frac{C-3a}{4}\right)^{2}=-\frac{5}{4}a^{2}+\frac{3}{2}Ca-\frac{C^{2}}{4}$. This is a parabola in $a$ with a maximum at $a_{0}=\frac{3C}{5}$. This proves the Proposition in the case $C=5k,k\in\mathbb{Z}$. In the case $C=5k+1,k\in\mathbb{Z}$ $a_{0}=\frac{3C}{5}$ lies in the interval $(3k,3k+1)$. For the values of $a,b_{i}$ from Proposition 4.4 (2) we have $Q=5k^{2}+2k-1$. A calculation gives $f_{5k+1,3k-1}=5k^{2}+2k-3$. Hence for all $a\leq 3k-1$ $Q\leq f_{5k+1,a}\leq f_{5k+1,3k-1}<5k^{2}+2k-1$. Also we find $f_{5k+1,3k+2}=5k^{2}+2k-\frac{9}{4}$, hence for all $a\geq 3k+2$ $Q\leq 5k^{2}+2k-\frac{9}{4}<5k^{2}+2k-1$. Hence $a\in\{3k,3k+1\}$. This proves the case (2). The cases (3-5) can be proven completely analogously. $\square$
@@
    $0\to\mathcal{O}_{X}(-1)^{\oplus 2}\to E\to\mathcal{O}_{S}(D)\to 0.$ (20)
  2. 2.
@@
Proof. We can calculate that the radius $\rho_{Q}$ of the numerical wall $Q_{\alpha,\beta}(E)=0$ satisfies the inequality
@@
Bounding each term, we see that $\frac{\rho_{Q}}{\overline{\Delta}_{H}(E)}>\frac{1}{300}$, hence we can apply Proposition 2.4 (1) to get the statement of our lemma. $\square$
Lemma 4.7
@@
Proof. Follows along the same lines as the proof of [23, Lemma 3.9]. Note that in our case
@@
Since the step by which $\mathrm{ch}_{3}(E)$ can change for a fixed $\mathrm{ch}_{\leq 2}(E)$ is equal to $\frac{H^{3}}{10}$, the bound proven in [23, Lemma 3.9] suffices to prove our lemma. $\square$
Список литературы
  • [1] E. Arrondo, L. Costa, ‘‘Vector bundles on Fano threefolds without intermediate cohomology’’, Comm. Algebra, 28:8 (2000), 3899–3911.
  • [2] A. Bayer, E. Macrм, P. Stellari, ‘‘The space of stability conditions on abelian threefolds, and on some Calabi-Yau threefolds“, Invent. math., 206 (2016), 869–933.
  • [3] A. Bayer, E. Macrм, Yu. Toda, ‘‘Bridgeland Stability conditions on threefolds I: Bogomolov-Gieseker type inequalities’’, J. Algebraic Geom., 23 (2014), 117–163.
  • [4] M. Bernardara, E. Macrм, B. Schmidt, X. Zhao, ‘‘Bridgeland Stability Conditions on Fano Threefolds’’, Йpijournal de Gйomйtrie Algйbrique, 1 (2017).
  • [5] T. Bridgeland, ‘‘Stability conditions on $K3$ surfaces’’, Duke Math. J., 141:2 (2008), 241–291.
  • [6] L. Ein, R. Hartshorne, H. Vogelaar, ‘‘Restriction Theorems for Stable Rank 3 Vector Bundles on $\mathbb{P}^{n}$’’, Math. Ann., 259 (1982), 541–569.
  • [7] D. Faenzi, ‘‘Bundles over the Fano threefold $V_{5}$’’, Comm. Algebra, 33:9 (2005), 3061–3080.
  • [8] D. Faenzi, ‘‘Even and odd instanton bundles on Fano threefolds of Picard number one’’, manuscripta math., 144 (2014), 199–239.
  • [9] W. Fulton. Intersection Theory, Springer New York, NY, 1998 (2nd edition), xiii+470 pp.
  • [10] R. Hartshorne, ‘‘Stable reflexive sheaves’’, Math. Ann., 254 (1980), 121–176.
  • [11] В. А. Исковских, Лекции по трехмерным алгебраическим многообразиям: Многообразия Фано, М.: Изд-во Моск. ун-та, 1988, 164 с.
  • [12] A. Kuznetsov, ‘‘Instanton bundles on Fano threefolds’’, Cent. Eur. J. Math., 10:4 (2012), 1198–1231.
  • [13] K.-S. Lee, K.-D. Park. ‘‘Moduli spaces of Ulrich bundles on the Fano 3-fold V5’’, Journal of Algebra, 574 (2021), 262–277.
  • [14] Ch. Li, ‘‘Stability conditions on Fano threefolds of Picard number 1’’, J. Eur. Math. Soc., 21:3 (2019), 709–726.
  • [15] E. Macrм, ‘‘A generalized Bogomolov-Gieseker inequality for the three-dimensional projective space’’, Algebra Number Theory, 8:1 (2014), 173–190.
  • [16] E. Macrм, ‘‘Stability conditions on curves’’, Mathematical Research Letters, 14 (2007), 657–672.
  • [17] E. Macrм, B. Schmidt. ‘‘Derived categories and the genus of space curves’’, Algebraic Geometry 7:2 (2020), 153–191.
  • [18] A. Nordskova, M. Van den Bergh. ‘‘Full exceptional collections on Fano threefolds and groups generated by spherical twists on K3 surfaces’’, arXiv:2412.06023v3 [math.AG].
  • [19] Ch. Okonek, M. Schneider, H. Spindler, Vector bundles on complex projective spaces, Progr. Math., vol. 3, Birkhaäuser, Boston, Mass. 1980, vii+389 pp.
  • [20] D. O. Orlov, “Exceptional set of vector bundles on the variety V5”, Vestnik Moskov. Univ., Ser. 1. Mat. Mekh., no. 5 (1991), 69–71.
  • [21] X. Qin, ‘‘Compactification of the moduli space of minimal instantons on the Fano 3-fold V5’’, Journal of Pure and Applied Algebra, 225:3, (2021), 106526.
  • [22] B. Schmidt, ‘‘A generalized Bogomolov-Gieseker inequality for the smooth quadric threefold’’, Bull. Lond. Math. Soc., 46:5 (2014), 915–923.
  • [23] B. Schmidt, ‘‘Rank two sheaves with maximal third Chern character in three-dimensional projective space’’, arXiv:1811.11951 [math.AG] (published in Matematica Contemporanea, 47 (2018), 228–270).
  • [24] B. Schmidt, ‘‘Bridgeland Stability on Threefolds – Some Wall Crossings’’, arXiv:1509.04608v2 [math.AG] (published in J. Algebraic Geom. 29 (2020), 247-283).
  • [25] B. Schmidt, ‘‘Sheaves of low rank in three-dimensional projective space’’, European Journal of Mathematics 9 (2023).
  • [26] V. do Valle Pretti, ‘‘Determinants, Even Instantons and Bridgeland Stability’’, Bull. Braz. Math. Soc., New Series 56:9 (2025).
  • [27] D. A. Vasil’ev, A. S. Tikhomirov, “Moduli of rank 2 semistable sheaves on rational Fano threefolds of the main series”, Sb. Math., 215:10 (2024), 1269–1320.
Danil A. Vassiliev
National Research University Higher School of Economics, Russian Federation
AG Laboratory, HSE, 6 Usacheva str., Moscow, Russia, 119048
E-mail: danneks@yandex.ru